OpenStudy (ezraweise):

Solve. 3y + 15 > –6 A. y > 3 B. y > –7 C. y > 13 D. y > –17

OpenStudy (anonymous):

change the original equation so it has only one y

OpenStudy (anonymous):

on the left side

OpenStudy (ezraweise):

?

OpenStudy (anonymous):

you want to compare the equation 3y + 15 > –6 with the answers. the answers all have one y on the left. So, apply equivalent operations left and right of the equation 3y + 15 > –6 until you have the same form

OpenStudy (ezraweise):

How? i am confused.

OpenStudy (anonymous):

you can always do the same operations on both sides of an equation, and the equation is still correct. 3y + 15 > –6 what operation would you want to do on the left side to get y alone?

OpenStudy (anonymous):

C

OpenStudy (ezraweise):

ThankS

OpenStudy (radar):

The above series of post is a good example where the desire for an answer exceeds the desire to learn. @phant_seea was in the process of teaching @ezraweise how to solve the given inequality when @gdbug interrupted providing an answer which was immediately given "thanks" for and the student moved on. Regardless that the answer was wrong, and the process on how to actually solve the problem was terminated. Here at Open Study, it is hoped some learning will take place. The events that occurred above prevented anything good happening. I am not pointing my finger at anybody, but if you review what happened during this problem, you can draw your own conclusion where the fault lies.

OpenStudy (anonymous):

first sold...3x+15=-6 3x=-6 -15=-21 x=-21/3 x=-7 then x>-7

OpenStudy (radar):

Too bad this step by step procedure did not get posted in time.

OpenStudy (paulodomingos):

\[3y+15>-6\] \[3y+15-15>-6-15\] \[3y >-21\] \[y >\frac{ -21 }{ 3 }\] \[y >-7\]